OpenStudy (anonymous):

find the polynomial function with -1/2, -2, 3

OpenStudy (jdoe0001):

assuming those are the roots? \({\bf -\cfrac{1}{2}, -2, 3\implies x=\{-\cfrac{1}{2}, -2, 3\}}\\ \quad \\ \begin{array}{llll} x=-\cfrac{1}{2}\implies x+\cfrac{1}{2}=0\implies &\left(x+\cfrac{1}{2}\right)=0\\ \quad \\ x=-2\implies =0\implies &(x+2)=0\\ \quad \\ x=3\implies x-3=0\implies &(x-3)=0 \end{array}\\ \quad \\\implies \bf \left(x+\cfrac{1}{2}\right)(x+2)(x-3)=0 \)

OpenStudy (jdoe0001):

hmmm \({\bf -\cfrac{1}{2}, -2, 3\implies x=\{-\cfrac{1}{2}, -2, 3\}}\\ \quad \\ \begin{array}{llll} x=-\cfrac{1}{2}\implies x+\cfrac{1}{2}=0\implies &\left(x+\cfrac{1}{2}\right)=0\\ \quad \\ x=-2\implies x+2=0\implies &(x+2)=0\\ \quad \\ x=3\implies x-3=0\implies &(x-3)=0 \end{array}\\ \quad \\\implies \bf \left(x+\cfrac{1}{2}\right)(x+2)(x-3)=0\) so multiply the "roots" to get the polynomial

OpenStudy (anonymous):

thank you i really appreciate it

OpenStudy (jdoe0001):

yw